TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.BorisBellesia on 2006-12-08 - 17:54

60A converter tests - Friday 7th Dec.

Advancement of today:

  • AB-PO activities on cells C24-30_R7
  • 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
  • AB-PO main activity ended today;
  • Polarity tests on cells C24-26_R7

Monday activities:

  • AB-PO - tests on faulty cells: - C34_L8 - 1 converter in C32_R7 - 1 converter in C12_R7 - no problems for coactivities.
  • Polarity and conneciton in C28-32_L8 / C30-34_R7; no problems for coactivities
  • ELQA in C18-26_L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-08 - 10:27

Preparation for Cool Down RAT - Dec 8

PDF RAT_20061208.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x08x10x17/RAT_20061208.pdf

IMPORTANT: PLEASE COMMENT ON THE OPEN ISSUES IF YOU FIND THEM INCORRECT, INCOMPLETE OR NOT-PERFECTLY UPDATED.

AC non-conformity

  • 07.12.06 Non conformity in DFBMH, DFBAN and DFBXG current lead heaters (P.Chambouvet-J-C.Guillaume)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling

  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum -

Cryogenics

  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 08 Dec 2006 - 10:17

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-12-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login